OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, -7) and (6, -5)?

OpenStudy (saifoo.khan):

(-5+7)/(6+6)

OpenStudy (saifoo.khan):

DYS

OpenStudy (anonymous):

So

OpenStudy (anonymous):

=2/12=1/6

OpenStudy (saifoo.khan):

DYS = Do yourself. ^_~